OpenStudy (anonymous):

What is the product of 3 1/5 times 5/6 then the answer in simplest from?

OpenStudy (anonymous):

\[3 1/2 = \frac{ 7 }{ 2 }\] \[\frac{ 7 }{ 2 } \times \frac{ 5 }{ 6 }= \frac{ 35 }{ 12 }\] Its simply 7x5=32 & 2x6=12

OpenStudy (anonymous):

a. three and five over eleven b. three and one over six c. two and two over three d. two and one over four

OpenStudy (anonymous):

those are my options

OpenStudy (paki):

7*5 = 35 and 2*6 = 12.. .so simple form will be 35/12...

OpenStudy (anonymous):

What is the product of three and one over five times five over six in simplest form?

OpenStudy (anonymous):

thats the question

OpenStudy (paki):

are you talking about this.... (3)* (1/5) / 6.... ?

OpenStudy (anonymous):

No its like 31/5 the fraction times 5/6 and the answer to the question is the problems answer in simpilest form.

OpenStudy (paki):

(31/5)*(5/6)....? is this rite....?

OpenStudy (anonymous):

Yeah so before 1/5 3 is a whole number in the fraction :)

OpenStudy (paki):

ahhh.. math is not difficult... but its expressions are difficult :-)

OpenStudy (anonymous):

haha yeah my brain gets so confused haha

OpenStudy (paki):

my too :-)

OpenStudy (paki):

please just write your question in the "draw" button and then post...

OpenStudy (anonymous):